# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 4.5 Exhibiting a super-solution to the evolution equation of the Fourier transform
In order to determine super-solutions of (4.16) in the sense of Definition 6, that is large enough to estimate from above the solutions of (4.16) in $\widetilde{\mathcal{C}}_{\infty,1}$, and that at the same time have an interesting long-time behaviour, we will, once again, take inspiration from [17], and consider functions of the form $u_{p}(t,\xi)=\varphi(t)|\xi|^{p}$.
Contrary to the situation in [17], we will see that $u_{p}$ decay and converge to zero as $t\rightarrow+\infty$, as soon as $p>0$, which will enable us to deduce the long-time behaviour of the solutions of (4.16).
@@
The key observation is that the functions of the form $C|\xi|^{p}$ are eigenfunctions of the gain collision operator $Q^{+}$ defined in (B.17). In fact, we have
@@
$\left|\frac{1-r}{2}e_{1}+\frac{1+r}{2}\sigma^{\prime}\right|^{p}$ $=\left(\left|\frac{1-r}{2}e_{1}+\frac{1+r}{2}\sigma^{\prime}\right|^{2}\right)^{p/2}$
$=\left(\left(\frac{1-r}{2}\right)^{2}+\frac{(1-r)(1+r)}{2}\sigma^{\prime}\cdot e_{1}+\left(\frac{1+r}{2}\right)^{2}\right)^{p/2}$
$=\left(\frac{1-2r+r^{2}+1+2r+r^{2}}{4}+\frac{1-r^{2}}{2}\sigma^{\prime}_{1}\right)^{p/2}$
@@
hence
$Q^{+}\left[|\xi|^{p}\right]=\lambda(p)|\xi|^{p},$ (4.55)
where $\lambda(p)$ is defined in (4.51), and $Q^{+}$ denotes the gain term of the collision operator of (B.15).
@@
Consider now the derivative of $\lambda$ with respect to $p$, i.e.
@@
We observe now that $-1<\sigma_{1}<1$ for almost every $\sigma\in\mathbb{S}^{d-1}$, then for almost every $\sigma$ we have
@@
which implies
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}p}\lambda(p)<0.$ (4.61)
Therefore, we get that $\lambda(p)<1$, for all $p>0$. ∎
@@ 4.6 Conclusion of the long-time behaviour argument
Let $\mathcal{F}_{0}$, $\mathcal{G}_{0}$ be the Fourier transforms $f,g\in\mathscr{\mathbb{R}^{d}}$ respectively. Assume that $f,g$ have the same mass and finite first order moments, i.e.
$\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)$ $=\int_{\mathbb{R}^{d}}g(\mathop{}\!\mathrm{d}v);$
$\int_{\mathbb{R}^{d}}|v|f_{0}(\mathop{}\!\mathrm{d}v)$ $=\int_{\mathbb{R}^{d}}|v|g(\mathop{}\!\mathrm{d}v).$ (4.62)
@@
In particular, if $\mathcal{F}_{\infty}$ denotes the Fourier transform of the steady state $f_{\infty}$ obtained in 1.2, we have
@@
By our assumption $f_{0},g_{0}$ we can apply Proposition 4.2, which gives that
@@
Finally from Proposition 4.6 and the explicit expression of $u_{1}$, we conclude that $u_{1}$ converges to zero for every compact set $K\subseteq\mathbb{R}^{d}$ as $t\rightarrow+\infty$. ∎
Using now classical results (see for instance [36]) we can deduce, from the pointwise convergence of the Fourier transforms $\mathcal{F}(t,\cdot)$ of the solutions $f(t)$ to (1.11), the weak convergence of the measures $f(t)$.
@@
Let $f_{0},g_{0}\in\mathscr{M}(\mathbb{R}^{d})$ with the same mass and finite first order moments:
$\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)$ $=\int_{\mathbb{R}^{d}}g(\mathop{}\!\mathrm{d}v);$
$\int_{\mathbb{R}^{d}}|v|f_{0}(\mathop{}\!\mathrm{d}v)$ $=\int_{\mathbb{R}^{d}}|v|g(\mathop{}\!\mathrm{d}v).$ (4.67)
Then, the two solutions $f,g$ of (1.11) given by Theorem 1.1 with respective initial data $f_{0}$ and $g_{0}$, are weakly converging to 00 as $t\rightarrow+\infty$.
@@ 5 Long-time behaviour of the moments for general homogeneity $\mu\geq 0$
In this section we will consider the evolution, and in particular the long-time behaviour, of the moments of the homogeneized version (1.1) of (1.2). We recall that the rehomogeneized homogeneous, inelastic linear Boltzmann equation, for Maxwell molecules, with a gravity field, is:
$\partial_{t}f(t,v)$ $+a\cdot\partial_{v}f(t,v)=T^{\mu}(t)\Bigg{[}\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 0.7113pt^{\prime}N\cdot\omega|)f(t,^{\prime}v)\mathop{}\!\mathrm{d}\omega-f(t,v)\Bigg{]},$
@@
As in Section 3.1 it is possible to derive the evolution equation for the moments of a solution of (1.1). Considering the weak formulation of the equation, and choosing $\{1,v,|v|^{2}\}$ as test functions we obtain the evolution equations of the moments of order zero, one and two. These evolution equations are explicitly written in the following proposition.
@@
Let $d=2,3$ and let the restitution coefficient $r$ be in ]0,1[]0,1[] 0 , 1 [. Let $\mu>0$ be a strictly positive real number. Let the angular collision kernel $b$ satisfy Assumption 1 and suppose that $f$ is a solution of (1.1). Then we have
@@
$\frac{\mathop{}\!\mathrm{d}T}{\mathop{}\!\mathrm{d}t}$ $=a\cdot M_{1}-\frac{C_{d}}{2}(1-r^{2})T^{\mu+1},$ (5.3)
where $C_{d}$ is constant depending only on $d$ for which we have
@@
The system of equations obtained in Proposition 5.1 is closed. Nevertheless, the non-linear structure of the system in the case $\mu>0$ prevents to solve it explicitly. Nevertheless, it is possible to obtain detailed information on the behaviour of the solutions to (5.2)-(5.3) with a qualitative study of the respective ODEs. Notice that independently on $\mu$ we always have $M_{0}(t)=M(0)$. Since the case $\mu=0$ was already addressed in Section 3.1, we will focus on the case $\mu\neq 0$.
@@
Denoting the temperature by $y$, we will obtain a two-dimensional system of ODEs of the form:
$\left\{\begin{array}[]{ccc}\vspace{2mm}\displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}x&=&f(x,y),\\ \displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}y&=&g(x,y),\end{array}\right.$ (5.8)
with $f,g:\mathbb{R}^{2}\rightarrow\mathbb{R}$. Namely, from (5.2)-(5.3), we deduce the following system of evolution equations for $x(t)$ and $y(t)$:
$\left\{\begin{array}[]{rcl}\vspace{2mm}\displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}x&=&|a|^{2}M_{0}-C_{d}(1+r)xy^{\mu},\\ \displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}y&=&2x-C_{d}(1-r^{2})y^{\mu+1}.\end{array}\right.$ (5.11)
@@
Nevertheless, another important constraint has to be kept in mind, since the first and second moments are not independent, because they are both linked to the distribution function $f$. Since we have:
$M_{1}(t)=\int_{\mathbb{R}^{d}}vf(t,v)\mathop{}\!\mathrm{d}v,$ (5.13)
we deduce that:
@@
As a consequence, the relevant region in the phase space of the ODE system on $(x,y)$ is the hypergraph of the function ${y=\frac{x^{2}}{|a|M_{0}}}$. We observe in particular that points of the form $(x,0)$, with $x<0$, do not belong to such a region, ruling out the only possibility to produce solutions with a temperature switching from positive, physical values, to negative values.
@@
Then, the region:
=$\mathcal{R}=\{(x,y)\in\mathbb{R}^{2}\ /\ |a|M_{0}y\geq x^{2}\}$ (5.15)
@@
At any point of coordinates $(x,y)$ of the boundary $x^{2}=|a|^{2}M_{0}y$ of the region $\mathcal{R}$, a normal, ingoing vector $n(x,y)$ to the boundary at this point is given by:
@@
Computing the scalar product between the ingoing normal vector $n$ and the time derivative of the integral curve of the system (5.11) through the point $(x,y)$ provides:
$n(x,y)\cdot\left(x^{\prime}(t),y^{\prime}(t)\right)$ $=\left(-2x,|a|^{2}M_{0}\right)\cdot\left(x^{\prime}(t),y^{\prime}(t)\right)$
$=-2x\left(|a|^{2}M_{0}-C_{d}(1+r)xy^{\mu}\right)+|a|^{2}M_{0}\left(2x-C_{d}(1-r^{2})y^{\mu+1}\right)$
$=2C_{d}(1+r)x^{2}y^{\mu}-C_{d}(1-r^{2})|a|^{2}M_{0}y^{\mu+1}$
$=C_{d}|a|^{2}M_{0}y^{\mu+1}\left[2(1+r)-(1-r^{2})\right]$
@@
We observe, on the one hand, that we can always assume $M_{0}>0$ (if not, the situation is trivial, for $f$ would describe void), and on the other hand that $a>0$. In the case $a=0$, that is, in the case of a Lorentz gas of Maxwell molecules, without any acceleration field, the projection $x$ we introduced is zero by definition, and the evolution equation of the temperature becomes the simple one-dimensional ODE:
@@
Therefore, in the “non trivial” case $a\neq 0$, $M_{0}\neq 0$, we find that $n(x,y)\cdot(x^{\prime},y^{\prime})>0$, proving that the hypergraph of $x^{2}=|a|^{2}M_{0}y$ is an invariant region under the flow of the ODE system (5.11). In other words, no solution of the ODE system starting in the hypergraph of $x^{2}=|a|^{2}M_{0}y$ can leave this region ∎
@@
Therefore, we have:
$2x_{\infty}^{2}$ $=C_{d}(1-r^{2})x_{\infty}y_{\infty}^{\mu}\cdot y_{\infty}$
$=C_{d}(1-r^{2})\frac{|a|^{2}M_{0}}{C_{d}(1+r)}y_{\infty},$ (5.20)
that is
$x_{\infty}^{2}=\frac{(1-r)}{2}|a|^{2}M_{0}y_{\infty},$ (5.21)
so that
@@ 5.2 Long time behaviour of the moments
If a solution $(x(t),y(t))$ satisfies $x(t_{0})=0$ for some time $t_{0}$, then $x^{\prime}(t_{0})>0$, so such a solution would be outside at $t_{0}-h$ for $h>0$ small enough. If $y(t_{0})=0$ and $x(t_{0})>0$ then $y^{\prime}(t_{0})>0$, and the same argument applies. If finally $x(t_{0})=y(t_{0})=0$, then the solution lies on the boundary of $\mathcal{R}$ at time $t_{0}$, which is an invariant region under the flow of (5.11) by Proposition 5.2.
We now prove that all the solutions enter eventually in $\mathcal{R}\cap\{(x,y)\in\mathbb{R}^{2}\ /\ x\geq 0,y\geq 0\}$. Let us assume that a solution $(x(t),y(t))$ of the system (5.11) remains for all positive time in the quadrant $x\leq 0,y\geq 0$. Since this solution remains in $\mathcal{R}$ by Proposition 5.2, we observe that $x^{\prime}\geq 0$, $y^{\prime}\leq 0$, so that the coordinate $x$ being bounded from above and the coordinate $y$ being bounded from below, the solution is globally defined, and its integral curve should converge towards a limit, which has to be an equilibrium, but there is no such an equilibrium in $\mathcal{R}$ intersected with the quadrant $x\leq 0,y\geq 0$, so all the solutions enter the region $x\geq 0,y\geq 0$ in finite time. ∎
Therefore, we need only to study phase portrait of the ODE system (5.11) in the first quadrant intersected with the invariant region $\mathcal{R}$, as pictured in Figure 1.
@@
The vertical isocline, characterized by the points $(x,y)$ of the phase space for which the integral curve through this point satisfies $x^{\prime}=0$, is the curve (represented in blue in Figure 1) with equation:
@@
In the same way, the horizontal isocline (such that $y^{\prime}=0$) is the curve (represented in red in Figure 1) of equation:
@@
We observe now that if a solution enters in one of the regions $R_{i}$, and remains in this region for any positive time, then the solution is global, and it converges towards the only equilibrium $(x_{\infty},y_{\infty})$. Let us present the arguments that allow to reach such a result, for the particular case of the region $R_{1}$. The cases of the three other regions can be studied in the exact same manner. So, we assume that we have a solution that remains in $R_{1}$ for any positive time of its interval of definition. In $R_{1}$, we have $x^{\prime}\geq 0$ and $y^{\prime}\geq 0$. In particular, we have $y(t)\geq y(0)$ for any $t\geq 0$ of the time interval of definition. The intersection between $R_{1}$ and $y\geq y(0)$ defines a bounded region of the phase space, we deduce therefore that such a solution has to be global. The coordinates $x$ and $y$ of the solution being bounded from above and increasing, they converge towards a finite limit as $t\rightarrow+\infty$. Such a limit has therefore to be an equilibrium, because if not, one of the derivatives $x^{\prime}$ or $y^{\prime}$ converges towards a non zero limit, which in turn contradicts the boundedness of the solution.
@@
With the exact same arguments, we would obtain that the only way to leave $R_{2}$ is to enter $R_{3}$, the only way to leave $R_{3}$ is to enter $R_{4}$, and the only way to leave $R_{4}$ is to enter $R_{1}$. Therefore, a solution that leaves any of the regions $R_{i}$ in finite time travels the regions according to the order $R_{1}\rightarrow R_{2}\rightarrow R_{3}\rightarrow R_{4}\rightarrow R_{1}\rightarrow\dots$, forever. Without loss of generality, we consider a solution initially starting from $\left(x_{0},y_{0}\right)\in R_{4}$. We can now deduce the following.
@@
  • The solution enters $R_{1}$ at a point $(x_{R_{1}},y_{R_{1}})$ such that $x_{R_{1}}\geq x_{1}$ and $y_{R_{1}}\geq y_{1}$. But now in $R_{1}$, we have $x^{\prime}\geq 0$ and $y^{\prime}\geq 0$. Therefore, in $R_{1}$ the solution remains in the region $y\geq y_{R_{1}}\geq y_{1}$. It implies that the solution remains in the region $x\leq x_{2}$, with:
[…]
  • The solution enter then $R_{2}$ at a point $(x_{R_{2}},y_{R_{2}})$, with $x_{R_{2}}\leq x_{2}$. In $R_{2}$, since $x^{\prime}\leq 0$, the solution remains in the region $x\leq x_{2}$. As a consequence, the solution remains in the region $y\leq y_{2}$, with:
@@
  • When the solution enters afterwards in $R_{3}$, it does at a point $(x_{R_{3}},y_{R_{3}})$ with $y_{R_{3}}\leq y_{2}$. In $R_{3}$ we have $y^{\prime}\leq 0$, so that the solution remains in the region $y\leq y_{2}$. Therefore, the solution remains in the region $x\geq x_{3}$ before it leaves $R_{3}$, with:
[…]
In the end, we obtained that when a solution starting from $(x_{0},y_{0})$ in $R_{1}$ leaves this region $R_{1}$, it re-enters this region $R_{1}$ for the first time at a point $(x_{R_{1}},y_{R_{1}})$ with:
@@
But by definition of the different values $x_{i}$, $y_{i}$, we have:

### v2

@@ 4.5 Exhibiting a super-solution to the evolution equation of the Fourier transform
Now, we determine super-solutions of (4.14) in the sense of Definition 6, that are large enough to estimate from above the solutions of (4.14) in $X$. To ensure in addition that the super-solutions have an interesting long-time behaviour, we will consider functions of the form $u_{p}(t,\xi)=\varphi(t)|\xi|^{p}$, as in [15].
Contrary to the situation in [15], we will see that $u_{p}$ decays and converges to zero as $t\rightarrow+\infty$, as soon as $p>0$, which will enable us to deduce the long-time behaviour of the solutions of (4.14).
@@
The key observation is that the functions of the form $C|\xi|^{p}$ are eigenfunctions of the gain part $Q^{+}$ of the collision operator defined in (A.14). In fact, we have
@@
$\left|\frac{1-r}{2}e_{1}+\frac{1+r}{2}\sigma^{\prime}\right|^{p}$ $=\left(\left(\frac{1-r}{2}\right)^{2}+\frac{(1-r)(1+r)}{2}\sigma^{\prime}\cdot e_{1}+\left(\frac{1+r}{2}\right)^{2}\right)^{p/2}$
@@
hence $Q^{+}\left[|\xi|^{p}\right]=\lambda(p)|\xi|^{p}$ where $\lambda(p)$ is defined in (4.44), and $Q^{+}$ denotes the gain term of the collision operator of (A.12).
@@
We consider now the derivative of $\lambda$ with respect to $p$, i.e.
@@
We observe now that $1<\sigma_{1}<1$ for almost every $\sigma\in\mathbb{S}^{d-1}$, then for almost every $\sigma$ we have
@@
which implies ${\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}p}\lambda(p)<0}$. Therefore, we get that $\lambda(p)<1$, for all $p>0$. ∎
@@ 4.6 Conclusion of the long-time behaviour argument
Let $\mathcal{F}_{0}$, $\mathcal{G}_{0}$ be the Fourier transforms of two probability measures $f_{0},g_{0}\in\mathcal{M}_{+}(\mathbb{R}^{d})$ respectively. Assume in addition that $f_{0},g_{0}$ have finite first order moments. In other words, we assume:
$\hskip 34.1433pt\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)=\int_{\mathbb{R}^{d}}g_{0}(\mathop{}\!\mathrm{d}v)=1,$
$\int_{\mathbb{R}^{d}}|v|f_{0}(\mathop{}\!\mathrm{d}v)<+\infty,\hskip 8.53581pt\int_{\mathbb{R}^{d}}|v|g_{0}(\mathop{}\!\mathrm{d}v)<+\infty.$ (4.53)
@@
In particular, if $\mathcal{F}_{\infty}$ denotes the Fourier transform of the steady state $f_{\infty}$ obtained in Theorem 1.2, we have
@@
By our assumption on $f_{0},g_{0}$ we can apply Proposition 4.2, which gives that
@@
Finally from Proposition 4.6 and the explicit expression of $u_{1}$, we conclude that $u_{1}$ converges to zero on every compact set $K\subseteq\mathbb{R}^{d}$ as $t\rightarrow+\infty$. ∎
Using now classical results (see for instance [30]) we can deduce, from the pointwise convergence of the Fourier transforms $\mathcal{F}(t,\cdot)$ of the solutions $f(t)$ to (1.9), the weak convergence of the measures $\big(f(t)\big)_{t}$.
@@
Let $f_{0},g_{0}\in\mathscr{M}(\mathbb{R}^{d})$ be two probability measures with finite first order moments:
$\hskip 34.1433pt\int_{\mathbb{R}^{d}}f_{0}(\mathop{}\!\mathrm{d}v)=\int_{\mathbb{R}^{d}}g_{0}(\mathop{}\!\mathrm{d}v);$
$\int_{\mathbb{R}^{d}}|v|f_{0}(\mathop{}\!\mathrm{d}v)<+\infty,\hskip 8.53581pt\int_{\mathbb{R}^{d}}|v|g_{0}(\mathop{}\!\mathrm{d}v)<+\infty.$ (4.58)
Then, the two solutions $f,g$ of (1.9) given by Theorem 1.1 with respective initial data $f_{0}$ and $g_{0}$, are weakly converging to 0 as $t\rightarrow+\infty$.
@@ 5 Long-time behaviour of the moments for general homogeneity $\mu\geq 0$
In this section we will consider the evolution, and in particular the long-time behaviour, of the moments of the homogeneized version (1.1) of (1.2). We recall that the rehomogeneized homogeneous, inelastic linear Boltzmann equation, for Maxwell molecules, with a gravity field, is
$\partial_{t}f(t,v)$ $+a\cdot\partial_{v}f(t,v)=T^{\mu}(t)\Bigg[\int_{\mathbb{S}^{d-1}}\frac{1}{r}b(|\hskip 0.7113pt^{\prime}N\hskip-1.42262pt\cdot\omega|)f(t,^{\prime}\hskip-2.84526ptv)\mathop{}\!\mathrm{d}\omega-f(t,v)\Bigg],$
@@
As in Section 3.1, it is possible to derive the evolution equation for the moments of a solution of (1.1). Considering the weak formulation of the equation, and choosing $\{1,v,|v|^{2}\}$ as test functions we obtain the evolution equations of the moments of order zero, one and two. These evolution equations are explicitly written in the following proposition.
@@
Let $d=2,3$ and let the restitution coefficient $r$ be in ]0,1[]0,1[. Let $\mu>0$ be a strictly positive real number. Let the angular collision kernel $b$ satisfy Assumption 1 and suppose that $f$ is a solution of (1.1). Then we have
@@
$\frac{\mathop{}\!\mathrm{d}T}{\mathop{}\!\mathrm{d}t}\hskip 2.84526pt$ $=a\cdot M_{1}-\frac{C_{d}}{2}(1-r^{2})T^{\mu+1},$ (5.3)
where $C_{d}$ is a constant depending only on $d$ for which we have
@@
The system of equations obtained in Proposition 5.1 is closed. However, the non-linear structure of the system in the case $\mu>0$ prevents to solve it explicitly. Nevertheless, it is possible to obtain detailed information on the behaviour of the solutions to (5.2)-(5.3) with a qualitative study of these ODEs. Notice that independently on $\mu$ we always have $M_{0}(t)=M(0)$. Since the case $\mu=0$ was already addressed in Section 3.1, we will focus on the case $\mu\neq 0$.
@@
Denoting the temperature by $y$, we will obtain a two-dimensional system of ODEs of the form $\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}(x,y)=\big(f(x,y),g(x,y)\big)$ with $f,g:\mathbb{R}^{2}\rightarrow\mathbb{R}$. Namely, from (5.2)-(5.3), we deduce the following system of evolution equations for $x(t)$ and $y(t)$:
$\left\{\begin{array}[]{rcl}\vskip 5.69054pt\displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}x&=&|a|^{2}M_{0}-C_{d}(1+r)xy^{\mu},\\ \displaystyle{\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}}y&=&2x-C_{d}(1-r^{2})y^{\mu+1}.\end{array}\right$. (5.8)
@@
In addition, the first and second moments are not independent, because they are both linked to the distribution function $f$. Indeed, we have
@@
As a consequence, the relevant region in the phase space of the ODE system on $(x,y)$ is the hypergraph of the function ${y=\frac{x^{2}}{|a|M_{0}}}$. We observe in particular that points of the form $(x,0)$, with $x<0$, do not belong to such a region, ruling out the only possibility to produce solutions with a temperature switching from positive and physical values, to negative values.
@@
Let $a\in\mathbb{R}^{d}$ be a fixed vector, and let $C_{d}$ be as in (5.4). Let $\mu>0$ be a strictly positive real number. Then, the region:
$\mathcal{R}=\{(x,y)\in\mathbb{R}^{2}\ /\ |a|M_{0}y\geq x^{2}\}$ (5.11)
@@
At any point of coordinates $(x,y)$ of the boundary $x^{2}=|a|^{2}M_{0}y$ of the region $\mathcal{R}$, a normal, ingoing vector $n(x,y)$ to the boundary at this point is given by
@@
Computing the scalar product between the ingoing normal vector $n$ and the time derivative of the integral curve of the system (5.8) through the point $(x,y)$ provides
$n(x,y)\cdot\left(x^{\prime}(t),y^{\prime}(t)\right)$ $=-2x\left(|a|^{2}M_{0}-C_{d}(1+r)xy^{\mu}\right)+|a|^{2}M_{0}\left(2x-C_{d}(1-r^{2})y^{\mu+1}\right)$
@@
We observe, on the one hand, that we can always assume $M_{0}>0$, and on the other hand that $a>0$. In the case $a=0$, that is, in the case of a Lorentz gas of Maxwell molecules, without any acceleration field, the projection $x$ we introduced is zero by definition, and the evolution equation of the temperature becomes the simple one-dimensional ODE:
@@
Therefore, in the “non trivial” case $a\neq 0$, $M_{0}\neq 0$, we find that $n(x,y)\cdot(x^{\prime},y^{\prime})>0$, proving that the hypergraph of $x^{2}=|a|^{2}M_{0}y$ is an invariant region under the flow of the ODE system (5.8). In other words, no solution of the ODE system starting in the hypergraph of $x^{2}=|a|^{2}M_{0}y$ can leave this region. ∎
@@
Therefore, we have $2x_{\infty}^{2}=C_{d}(1-r^{2})x_{\infty}y_{\infty}^{\mu}\cdot y_{\infty}=C_{d}(1-r^{2})\displaystyle{\frac{|a|^{2}M_{0}}{C_{d}(1+r)}}y_{\infty}=(1-r)|a|^{2}M_{0}y_{\infty}$, so that
@@ 5.2 Long time behaviour of the moments
If a solution $(x(t),y(t))$ satisfies $x(t_{0})=0$ for some time $t_{0}$, then $x^{\prime}(t_{0})>0$, so such a solution would be outside at $t_{0}-h$ for $h>0$ small enough. We obtain therefore a contradiction by considering a solution starting inside $\mathcal{R}$ and by considering for $t_{0}$ the first time this solution leaves the first quadrant. If $y(t_{0})=0$ and $x(t_{0})>0$ then $y^{\prime}(t_{0})>0$, and the same argument applies. If finally $x(t_{0})=y(t_{0})=0$, then the solution lies on the boundary of $\mathcal{R}$ at time $t_{0}$, which is an invariant region under the flow of (5.8) by Proposition 5.2.
We now prove that all the solutions starting inside $\mathcal{R}$ enter eventually in $\mathcal{R}\cap\{(x,y)\in\mathbb{R}^{2}\ /\ x\geq 0,y\geq 0\}$. Let us assume that a solution $(x(t),y(t))$ of the system (5.8) remains for all positive time in the quadrant $x\leq 0,y\geq 0$. Since this solution remains in $\mathcal{R}$ by Proposition 5.2, we observe that $x^{\prime}\geq 0$, $y^{\prime}\leq 0$, so that the coordinate $x$ being bounded from above and the coordinate $y$ being bounded from below, the solution is globally defined, and its integral curve should converge towards a limit, which has to be an equilibrium, but there is no such an equilibrium in $\mathcal{R}$ intersected with the quadrant $x\leq 0,y\geq 0$, so all the solutions enter the region $x\geq 0,y\geq 0$ in finite time. The proof is now complete. ∎
Therefore, we need only to study the phase portrait of the ODE system (5.8) in the first quadrant intersected with the invariant region $\mathcal{R}$, as pictured in Figure 1.
@@
The vertical isocline, characterized by the points $(x,y)$ of the phase space for which the integral curve through this point satisfies $x^{\prime}=0$, is the curve (represented in blue in Figure 1) of equation
@@
In the same way, the horizontal isocline (such that $y^{\prime}=0$) is the curve (represented in red in Figure 1) of equation
@@
We observe now that if a solution enters in one of the regions $R_{i}$, and remains in this region for any positive time, then the solution is global, and it converges towards the only equilibrium $(x_{\infty},y_{\infty})$. Let us present the arguments that allow to reach such a result, for the particular case of the region $R_{1}$. The cases of the three other regions can be studied in the exact same manner. So, we assume that we have a solution that remains in $R_{1}$ for any positive time of its interval of definition. In $R_{1}$, we have $x^{\prime}\geq 0$ and $y^{\prime}\geq 0$. In particular, we have $y(t)\geq y(0)$ for any $t\geq 0$ of the time interval of definition. The intersection between $R_{1}$ and $y\geq y(0)$ defines a bounded region of the phase space, we deduce therefore that such a solution has to be global. The coordinates $x$ and $y$ of the solution being bounded from above and increasing, they both converge towards finite limits as $t\rightarrow+\infty$. Such a limit has therefore to be an equilibrium, because if not, one of the derivatives $x^{\prime}$ or $y^{\prime}$ converges towards a non zero limit, which in turn contradicts the boundedness of the solution.
@@
With the exact same arguments, we would obtain that the only way to leave $R_{2}$ is to enter $R_{3}$, the only way to leave $R_{3}$ is to enter $R_{4}$, and the only way to leave $R_{4}$ is to enter $R_{1}$. Therefore, a solution that leaves any of the regions $R_{i}$ in finite time travels the regions according to the cyclic order $R_{1}\rightarrow R_{2}\rightarrow R_{3}\rightarrow R_{4}\rightarrow R_{1}\rightarrow\dots$, forever. Without loss of generality, we consider a solution initially starting from $\left(x_{0},y_{0}\right)\in R_{4}$. We can now deduce the following.
@@
  • The solution enters $R_{1}$ at a point $(x_{R_{1}},y_{R_{1}})$ such that $x_{R_{1}}\geq x_{1}$ and $y_{R_{1}}\geq y_{1}$. But now in $R_{1}$, we have $x^{\prime}\geq 0$ and $y^{\prime}\geq 0$. Therefore, in $R_{1}$ the solution remains in the region $y\geq y_{R_{1}}\geq y_{1}$. It implies that the solution remains in the region $x\leq x_{2}$, with
    $y_{1}=bx_{2}^{-1/\mu}\hskip 8.53581pt\text{that is}\hskip 8.53581ptx_{2}=\left(y_{1}/b\right)^{-\mu}.$ (5.21)
  • The solution enter then $R_{2}$ at a point $(x_{R_{2}},y_{R_{2}})$, with $x_{R_{2}}\leq x_{2}$. In $R_{2}$, since $x^{\prime}\leq 0$, the solution remains in the region $x\leq x_{2}$. As a consequence, the solution remains in the region $y\leq y_{2}$, with
@@
  • When the solution enters afterwards in $R_{3}$, it does at a point $(x_{R_{3}},y_{R_{3}})$ with $y_{R_{3}}\leq y_{2}$. In $R_{3}$ we have $y^{\prime}\leq 0$, so that the solution remains in the region $y\leq y_{2}$. Therefore, the solution remains in the region $x\geq x_{3}$ before it leaves $R_{3}$, with
    $y_{2}=bx_{3}^{-1/\mu}\hskip 8.53581pt\text{that is}\hskip 8.53581ptx_{3}=(y_{2}/b)^{-\mu}.$ (5.23)
In the end, we obtained that when a solution starting from $(x_{0},y_{0})$ in $R_{1}$ leaves this region $R_{1}$, it re-enters this region $R_{1}$ for the first time at a point $(x_{R_{1}},y_{R_{1}})$ with
@@
But by definition of the different values $x_{i}$, $y_{i}$, we have